Filter Corp. has a project available with the following cash flows: Year Cash Flow 0 −$16,000 1 5800​ 2 7100​ 3 6100​ 4 4900​ What is the project's IRR? Group of answer choices 18.84% 20.93% 21.98% 19.62% 20.41%
Answer #1
N A B=A/((1+R)^N) R=0.1884
YEAR CASHFLOW Present Value
0 ($16,000) ($16,000)
1 $5,800 $4,880.66
2 $7,100 $5,027.58
3 $6,100 $3,634.80
4 $4,900 $2,456.96
SUM ($0)
IRR (Internal Rate of Return)=Discount rate at which the Net Present Value =Zero
Assume , R=Internal Rate of Return
.-16000+5800/(1+R)+7100/((1+R)^2)+6100/((1+R)^3+4900/((1+R)^4)=0
R=Internal Rate of Return =IRR= 18.84% (Using IRR function of excel over the CASH FLOW)
ANSWER: 0.1884
18.84% Clipboard Font Allahment Number Styles C14 X for EIRR(D4:08) А E F G H NA B=A/((1+R)^N) R=0.1884 YEAR CASHFLOW Present Value